hartnn (hartnn):

Vector Calculus Prove : (a is constant vector.) \[\nabla \times [\dfrac{a\times \bar r}{r^n}] = \dfrac{2-n}{r^n}a + \dfrac{n(a.\bar r)\bar r}{r^{n+2}}\] So i used (any other method?) the identity \(\nabla \times \left( \mathbf{v \times F} \right) = \left[ \left( \mathbf{ \nabla \cdot F } \right) + \mathbf{F \cdot \nabla} \right] \mathbf{v}- \left[ \left( \mathbf{ \nabla \cdot v } \right) + \mathbf{v \cdot \nabla} \right] \mathbf{F}\)

hartnn (hartnn):

plugged in \(\nabla .a = 0 \\ \nabla .\bar r =3\)

hartnn (hartnn):

in \(\nabla \times \left( \mathbf{a \times \bar r } \right) = \left[ \left( \mathbf{ \nabla \cdot \bar r } \right) + \mathbf{\bar r \cdot \nabla} \right] \mathbf{a}- \left[ \left( \mathbf{ \nabla \cdot a } \right) + \mathbf{a \cdot \nabla} \right] \mathbf{\bar r}\)

hartnn (hartnn):

what next ?

hartnn (hartnn):

also if you see, i just removed r^n outside...is that legal ?

OpenStudy (dan815):

where did u get div a =0 and div r'=3 from

hartnn (hartnn):

oh, is that incorrect ?

OpenStudy (dan815):

no, jw

hartnn (hartnn):

\(\operatorname{div}\,\mathbf{F} = \nabla\cdot\mathbf{F} =\frac{\partial U}{\partial x} +\frac{\partial V}{\partial y} +\frac{\partial W}{\partial z }\)

OpenStudy (dan815):

ohh ok i see equating from that identity

hartnn (hartnn):

what does \(\bar r . \nabla\) even mean ?? and how di i calculate it ?

OpenStudy (dan815):

lol i was looking at that.. i was thought when it goes that way.. its just undefined

hartnn (hartnn):

it has to be a scalar, thats what i am sure of

OpenStudy (dan815):

taught* >_> and umm, what did u get for the expression for F

hartnn (hartnn):

F here is r = xi +yj +zk

OpenStudy (dan815):